Find the monthly payment that will yield the given future value. (Round your answer to the nearest cent.) $240,000 at 10 1/2 % interest for forty-two years; ordinary annuity
Answer #1

We will look at the solution in 3 ways using Excel, Financial calculator. Which is almost the same and similar process.

1. Excel

Input the data one by one as FV=240000, N=42 (42*12=504 as it is for monthly payments for 42 years), I or I/Y=10.50%(10.50%/12=0.00875 as it is annual rate convert it into monthly rate by dividing it by 12), Remember PV=0

Using Function PMT - Input same as below i a cell of excel

=PMT(0.00875,504,0,-240000,0) - This is how you input the data in a cell to calculate Monthly payment to be made to obtain 240000 at 10.50% for 42 years.

= You should get the answer as Rs.26.35 after doing the above method.

Coming to the financial calculater

Input the same - 240000(FV), 0.00875(I/Y), 504(N) - press (CPT) - Press (PMT) = 26.35 is the outcome.

Which is very much easy with the financial calculator.